A company manufactures three types of wooden chairs; the Kitui, the Goa, and the Santa Fe. To make a Kitui chair requires one hour of cutting time, 1.5 hours of assembly time, and one hour of finishing time. A Goa chair requires 1.5 hours of cutting time, 2.5 hours for assembly, and 2 hours for finishing.
The Santa Fe requires 1.5 hours of cutting, 3 hours of assembly, and 3 hours of finishing.If 41 hours of cutting time, 70 hours of assembly, and 58 hours of finishing time were used one week, how many of each type of chair were produced?

Answer by stanbon(48568)

A company manufactures three types of wooden chairs; the Kitui, the Goa, and the Santa Fe.
To make a Kitui chair requires one hour of cutting time, 1.5 hours of assembly time, and one hour of finishing time.
A Goa chair requires 1.5 hours of cutting time, 2.5 hours for assembly, and 2 hours for finishing.
The Santa Fe requires 1.5 hours of cutting, 3 hours of assembly, and 3 hours of finishing.
If 41 hours of cutting time,
70 hours of assembly,
and 58 hours of finishing time were used one week,
how many of each type of chair were produced?
------------------
Cutting: 1.5K+1.5G+1.5S=41
Asemble: 1.5K+2.5G+3S = 70
Finish: 1K +2G +3S = 58
-------------------
Solve the three equations:
I used matrices to get the solution:
K=0; G=24; S=3 1/3
Cheers,
Stan H.